# Strategic Informed Trading and the Value of Private Information

### v1

@@ Abstract.
We consider a market of risky financial assets where the participants are an informed trader, a mass of uniformed traders and noisy liquidity providers. We prove the existence of a market-clearing equilibrium when the insider internalizes her power to impact prices. In the price-impact equilibrium the insider strategically reveals a noisier (compared to when the insider takes prices as given) signal, and prices are less reactive to the publicly available information. In contrast to the related literature, we show that in the price-impact equilibrium, the insider’s ex-ante welfare monotonically increases in the signal precision. This clarifies when a trader with market power is motivated to both obtain and refine her private information. Furthermore, even though the uniformed traders act as price-takers, the effect of price impact is ex-ante welfare improving for them. By contrast, internalization of price impact may reduce insider ex-ante welfare. This happens provided the insider is sufficiently risk averse and the uninformed traders are sufficiently risk tolerant.
@@ Introduction
It is well-documented that large financial institutions possess the power to affect markets (see, for example, Koijen and Yogo (2019) and the related discussion in Rostek and Yoon (2023)). Compared to other traders, orders coming from large investors impact both transaction prices and volumes, and large investors are aware of this impact (see Rostek and Weretka (2015a)). Large financial institutions are additionally known to invest considerable capital to acquire information regarding the payoffs of the assets they trade (c.f. Kacperczyk and Pagnotta (2019)). In sum, it is natural to assume that large investors are both (1) aware of their impact on prices and (2) in possession of private information and hence are “insiders” compared to the other traders.
@@
That large investors are informed traders is not a secret. Indeed, the rest of the market knows large investors have private information (see Subrahmanyam (1991)), and as such, the uninformed traders’ account for the insider’s superior information, and expect that in equilibrium there is a partial transmission of private information. In both the price-taking (see Grossman and Stiglitz (1980)) and price-impact (see Kyle (1985)) cases, it is shown that in the presence of noise traders, the insider’s private signal is partially revealed to the uniformed traders through equilibrium prices, a mechanism that creates a market (or public) signal. Therefore, it is reasonable to assume that not only will the insider, who is also aware of her market power, strategically choose the signal she reveals to the market, but also that the uniformed traders recognise, and account for, this fact.
@@
Using the above as motivation, our goal is to study how the insider’s awareness of price impact affects equilibrium prices, information transmission and agent welfare. We work in the classic single period normal-CARA setting, seeking a linear price-impact equilibrium where a risk averse insider (endowed with private information) trades a bundle of risky assets with a mass of uniformed risk-averse traders, as well as liquidity providers (noise traders). We use this model to predict how the insider (partially) reveals her signal; how the uniformed agents correspondingly adjust their demands; and if the informational content within equilibrium prices is reduced compared to when all agents act as price takers.
We are also interested in agent welfare. Here, we wish to know if internalization of price impact is welfare improving for the insider and/or the uninformed agent, and especially if the insider’s welfare increases with her signal precision. The latter question arises because in the price-taking equilibrium of Grossman and Stiglitz (1980), insider welfare does not always increase with the signal precision. More precisely, if the insider receives the signal $G=X+Z_{I}$ where $X$ is the asset payoff and $Z_{I}$ is a noise term with precision $p_{I}$, then the map “$p_{I}\to\textrm{Insider Welfare}$” is not monotonically increasing. This is shown in comparison to the price-impact case in Figure 1 below, but qualitatively the map takes the two forms shown in the figure.
@@
This is problematic, because either it is never advantageous to obtain the signal (the dot-dashed line) or it is beneficial to obtain the signal, but not to refine it beyond a certain precision level (solid line). As presumably there is a cost (in terms of money and/or effort) to produce the signal, the former case suggests price taking equilibria with private signals are somewhat artificial, and the latter case requires the insider to estimate model parameters (which are very hard to estimate especially for other agents) to know if she should try and improve her signal.
Given this, we would like to know in what environment the insider can be assured that a better quality signal is always better for her, and evidently this is not the case when all agents take prices as given. Furthermore, lack of monotonicity was also shown in the recent Nezafat and Schroder (2023) where all agents account for price impact. However, in our model, by differentiating the insider on two fronts: private information and internalization of market power, we establish monotonicity with respect to the signal precision (Proposition 3.2), and avoid the paradoxical situations where the insider is not motivated to refine her signal, even if doing so does not incur additional costs.
@@
We adjust the single period CARA-normal setting of Grossman and Stiglitz (1980) [^1] Grossman and Stiglitz (1980) assumes all traders are price takers, an assumption ubiquitous in the literature on equilibrium under heterogeneous information. Indeed, it was assumed in the seminal papers of Grossman (1976), Grossman and Stiglitz (1980) and Hellwig (1980), and with the exception of the literature strand started by Kyle (1985); Back (1992), Rochet and Vila (1994) and Subrahmanyam (1991), price taking has remained the dominant assumption. by allowing the insider to internalize her price impact, while maintaining the presence of price-taking uniformed traders and liquidity providers [^2] Assuming the uninformed traders are price takers is realistic, as they represent a mass of small risk-averse traders who rationally optimize their positions, but do not have the power to move prices.. Following the related literature (e.g. Kyle (1985); Rochet and Vila (1994)), we study a linear impact equilibrium where the insider perceives the market price to be an affine function of the sum of her and the noise traders’ demand [^3] Linear price-impact is common in the literature (see among others Kyle (1989), Vayanos (1999), Vives (2011)), and the affine structure of price impact is also seen (though not perceived a-priori by the insider) in the price taking, or competitive, equilibrium (see (26) below)., with the affine coefficients endogenously determined through market clearing.
@@
Our first main result (Theorem 1.6) establishes existence of an affine price impact equilibrium. The affine coefficients (which are not the same as in the price-taking case), are governed by the unique positive root of a certain cubic equation, given in (18) [^4] This stands in contrast to Subrahmanyam (1991) where equilibria are governed by solutions to a quintic equation. The difference arises because, consistent with Grossman and Stiglitz (1980) (see also Rochet and Vila (1994)) we assume the insider, by viewing both her demand and the resultant price, is able to deduce the noise trader’s demand.. For the sake of comparison, we summarize the price-taking results of Grossman and Stiglitz (1980) in Proposition 1.7, and for both the price taking and price impact cases we establish equilibria absent private information by showing it coincides with the zero precision limit: see Proposition 1.11.
@@
Having established existence of equilibria we perform a comparison analysis. While analysis of information transmission and price reactivity yields expected conclusions, the study of indirect utilities (or welfare) results in surprising and novel outcomes that stand in sharp contrast to the literature. Relating to information transmission, we show (similarly to Kacperczyk et al. (2023)) that the insider’s price impact makes the market signal fuzzier than in the competitive setting (see Proposition 2.1). This is reasonable, as the insider has a motive to hide her signal when submitting her demand, and hence her trading lowers the precision of the signal revealed by the prices. Conversely, the uniformed traders recognize the insider reveals a muddied signal, and respond with a less elastic demand function, which in turn makes the equilibrium price less reactive to the public information (consistent with the adverse-selection concerns of Kyle (1989) and Lou and Rahi (2023)). These results are robust across model parameter values (in particular the uninformed traders and insider need not have the same risk aversion) and we conclude that by assuming the insider is a price taker, one is implicitly assuming the market gets a more precise signal, and prices are more reactive to public information, than if the insider had internalized price impact.
We then turn our attention to traders’ indirect utility, which (with a slight abuse of terminology) we refer to as “welfare”. Our main result is that when the insider internalizes price impact (and the uninformed agents do not), insider ex-ante welfare is monotonically increasing in the private signal precision. As discussed above, this result does not hold when either (i) all agents are price takers (e.g. Grossman and Stiglitz (1980)) or (ii) all agents internalize price impact (e.g. Nezafat and Schroder (2023)). Thus, in the reasonable setting of a strategic insider, price-taking uniformed traders and noisy liquidity providers, the meaningful statement “better information increases welfare” always holds, and there is always an incentive to refine the signal. Qualitatively, this result holds because the insider’s strategic trading increases the value of her private signal, and hence refining her signal is attractive.
Our second welfare result stems from a desire to isolate the effects of price impact internalization and private information on insider welfare. In other words, we ask: can one attribute any gain (or loss) in agent welfare to just price impact internalization or just asymmetric information? To answer this, we must turn on and off both the internalization of impact and private information (hence the need for the four equilibria established in Section 1). Our comparison analysis yields several interesting outcomes. First, we show that absent a private signal, internalizing price impact always improves insider welfare. While this result may be expected at the ex-ante level, remarkably it holds for any realization of the noise trader demand. Hence, under symmetric information, the equilibrium at which the insider acts strategically (solely due to her market power) yields a higher utility gain. We conclude that price impact internalization alone increases insider welfare.
The situation changes when we layer in the private signal. While typically (i.e. across the bulk of the parameter space) internalization of price impact yields higher insider welfare, insider welfare does not always increase. Indeed, when the insider is sufficiently risk averse, the uninformed traders are sufficiently risk tolerant and the variance of noise demand is sufficiently low [^5] Note that the variance of noise traders demand can be seen as a proxy for the size of their trading. This approach has been used by Kovalenkov and Vives (2014a) and more recently by Nezafat and Schroder (2023)., internalization of price impact actually reduces the insider’s welfare (the exact condition is given in equation (36) of Proposition 3.6).
This holds because having private information reduces the asset’s risk for the insider, which in turn increases her expected demand (in line with the empirical evidence of Kacperczyk and Pagnotta (2019)). On the other hand, internalization of price impact makes her “hide” her private signal, and this is expected to reduce her demand. This means that price impact and private signal have opposite expected effects on the insider’s demand. While the equilibrium price decreases due to price impact, the insider’s demand is lower under the presence of signal when she is also sufficiently risk averse. If, in addition, the uniformed traders are highly risk tolerant, their demand increases and the insider’s trade at market-clearing equilibrium is reduced even further. In other words, under a large deviation of the traders’ risk aversions, the effect of price impact is dominated by the that of heterogeneous information, the insider’s demand is reduced and her utility gain is lower. However, when the traders have the same risk aversion, the effect of price-impact internalization is always beneficial for the insider. We conclude that both asymmetric information and deviation of agent risk aversions may reduce insider welfare in the price-impact equilibrium. Further details are given in Section 4.
@@
The above results underscore an important (and clarifying) fact. As long as the insider internalizes her price impact, equilibrium prices cannot be driven to the corresponding price-taking equilibrium ones. This stems from the uniformed traders’ optimal demand. Although they are price takers, when determining their optimal demand, they do recognize that the insider internalizes her price impact. As mentioned above, price impact reduces public signal’s precision, and hence uniformed traders demand becomes less elastic than in the competitive equilibrium, and this is true for any realization of the public signal and level of prices. The lower sensitivity on the public signal alters the intercept point of their demand function (as it depends on public signal). Hence, in the price-impact equilibrium the insider trades against a different residual demand than in the price-taking equilibrium, which implies that price-taking equilibrium cannot be written as a special case of the price-impact one.
Finally, analysis of the uniformed traders’ welfare also yields quite interesting outcomes. First, traders are at the same side of trade in both the price-impact and price taking equilibria (e.g. in both equilibria, all agents either buy or sell off their initial endowments). As the insider’s price impact decreases the equilibrium prices, the uniformed traders satisfy their optimal demand at a discount, due to the insider’s price impact. This holds regardless the presence of private signal, which means that price impact is ex-ante beneficial for the uniformed traders with and without asymmetric information, as shown in Proposition 3.5. This allows us to conclude that for the bulk of the parameter space, aggregate welfare of both insider and uniformed traders increases due to the internalization of price impact, and this holds with and without the private signal.
@@
Our paper contributes to the on-going literature on price-impact equilibria under asymmetric information, on market participants’ welfare and on the informativeness of equilibrium prices.
@@
In our price-impact equilibrium, the insider does not act as price-taker. Usually, the price-taking assumption is made for tractability, as in its absence one must specify a price impact model, and depending upon the specification, it may be very difficult to establish equilibria. In the aforementioned Kyle (1985); Back (1992); Rochet and Vila (1994), the insider’s demand is combined with exogenous noise traders’ demand before being sent to a risk neutral market maker who prices in a competitive environment. In Subrahmanyam (1991), market makers are allowed to be risk averse while quoting prices to remain at utility indifference (as opposed the uninformed agent of Grossman and Stiglitz (1980) who can be thought of as a market maker who quotes utility-optimal prices), but the insider does not know the noise trader demand before submitting her order (as in Kyle (1985)). Regarding Subrahmanyam (1991), our analysis updates in two directions. First, by assuming the uninformed agent is a utility optimizer, and second, by allowing the insider to the identify the noise trader demand through the public equilibrium price (as in Rochet and Vila (1994)).
@@
In the spirit of the seminal work of Kyle (1989), several models on normal-CARA setting with price-impact and asymmetric information have been developed. For example, Vayanos (2001) and Rostek and Weretka (2015a) study dynamic thin markets with and without market makers respectively, Rostek and Weretka (2015b) emphasizes on the traders’ interdependent preferences and correlated private signals, while Malamud and Rostek (2017) and Anthropelos and Kardaras (2024) consider decentralized exchanges and restricted participation settings respectively. Also, Bergemann et al. (2021) studies a market of divisible goods where agents receive correlated signals and their demand affects the revealed signal at equilibrium, as in our model. In these works, as in Kyle (1989) and Vives (2011), strategic agents submit demand schedule forming a Nash equilibrium. In contrast to our model, they consider all non-noise agents to be strategic, with private information mostly appearing only on investors’ endowments. An extensive recent overview of this literature is provided in Rostek and Yoon (2023).
While it is an undoubted fact that large financial institutions invest to obtain private information, even when trading in markets that are not thin, theoretical studies that justify the positive relation between better information and higher gains from trading are scarce. We highlight that a key factor which always leads to this positive relationship is the insider’s market power in a markets with a mass of small uniformed traders and noisy liquidity providers. Under a competitive market setting, the fact that private information has positive value if an investor acts strategically was pointed out back in Hirshleifer (1971), while similar positive effect of private signal on traders’ welfare is shown in a competitive model with a continuum of traders in Morris and Shin (2002).
@@
Under non-competitive market settings, information acquisition has been relatively recently studied in Vives (2011), Rostek and Weretka (2012) and Vives (2014), where insiders with correlated noisy signals are considered. An extension of these papers to a two-stage model has been recently developed in Nezafat and Schroder (2023). Therein, at the first stage, one type of traders can choose the precision of the private signal that they will get before trading. This situation is comparable to our model when the rest of the non-noise traders are uniformed (no private signal). In contrast to our model however, both the insider and uniformed traders are assumed strategic, which essentially implies that the market is thin despite the presence of noise traders. Strategic uniformed traders, together with specific conditions on noise traders’ demand, lead to an equilibrium at which private information is welfare-deteriorating. We show that the existence of zero-information equilibria is not possible if the uniformed traders are price takers. In this case, the insider’s ex-ante welfare is always increasing with respect to signal’s precision, which means that the private information does have positive value for the insider. In addition, we show that under Pareto-allocated initial endowments, private information has positive value for uniformed traders too (even though they are assumed as price-takers).
On the other hand, Kacperczyk et al. (2023) considers strategic informed and price-takers uniformed traders as we do, where the role of initial endowments is highlighted for different types of informed traders. In Gong et al. (2022) the insider is assumed risk neutral and the role of uniformed traders is played by an ambiguous market maker with a quadratic objective. A linear equilibrium, similar to Kyle (1985), is derived where the price is underreacted to public signal, as in our case.
@@
Strategic agents and asymmetric information have been included in Lou and Rahi (2023), where in a non-competitive market (in line with Kyle (1989) and Rostek and Weretka (2015a)) traders receive different ex-ante random values of a single asset (and potentially different private signals too). In contrast to our model (and to Kyle (1989)) there are no liquidity providers, which essentially implies that price informativeness does not change due to price impact. As in our model, there are conditions that lead to higher ex-ante expected utility for the uniformed traders than the insider. We reach a similar conclusion, but without assuming that uniformed traders act strategically.
@@
Another channel on the valuation of private information that leads to insider’s lower gains from trading due to private information is the information sharing. For example, Goldstein et al. (2023) reach to this result in a novel model (based on Kyle (1989)) where informed traders share their private signals before trading (as in Indjejikian et al. (2014)).
@@
The rest of the paper is organized as follows. In Section 1 provide the model and establish existence of the equilibria under consideration. Section 2 develops quantitative analysis and qualitative discussion on information transmission and signal and price sensitivities. Section 3 focuses on welfare comparison at different equilibria and Section 4 is dedicated to equilibria structure regarding prices and risk allocation and concludes with a discussion on model’s predictions. An extension to a general multi asset model is provided in Appendix A and all proofs are given, starting in Appendix B.

### v2

@@ Abstract.
We consider a market of risky financial assets whose participants are an informed trader, a representative uninformed trader, and noisy liquidity providers. We prove the existence of a market-clearing equilibrium when the insider internalizes her power to impact prices, but the uninformed trader takes prices as given. Compared to the associated competitive economy, in equilibrium the insider strategically reveals a noisier signal, and prices are less reactive to publicly available information. Additionally, and in direct contrast to the related literature, in equilibrium the insider’s indirect utility monotonically increases in the signal precision. Therefore, the insider is motivated not only to obtain, but also to refine, her signal. Lastly, we show that compared to the competitive economy, the insider’s internalization of price impact is utility improving for the uninformed trader, but somewhat surprisingly may be utility decreasing for the insider herself. This utility reduction occurs provided the insider is sufficiently risk averse compared to the uninformed trader, and provided the signal is of sufficiently low quality.
@@ Introduction
It is well-documented that large financial institutions possess the power to affect markets (for example, see Koijen and Yogo (2019) and Rostek and Yoon (2024)). Compared to other traders, large investors’ orders impact both transaction prices and volumes, and these investors are aware of this impact (Rostek and Weretka (2015a)). Additionally, financial institutions invest considerable capital to acquire information regarding traded asset payoffs (Kacperczyk and Pagnotta (2019)). In sum, it is natural to assume large investors are both aware of their impact on prices and in possession of private information.
@@
However, it is not a secret that large investors are informed traders (Subrahmanyam (1991)). Indeed, other market participants, even if they lack private information themselves, both know and account for the large investors’ (“insiders”) superior information. This means that in equilibrium, one expects a partial transmission of private information. Indeed, this was shown concretely in both competitive economy and price impact models (starting with Grossman and Stiglitz (1980) and Kyle (1985) respectively), as therein the insider’s private signal about the risky assets’ terminal payoff is partially revealed to all market participants through equilibrium prices, a mechanism that creates a market (or public) signal. Therefore, it is also reasonable to assume uninformed traders know the insider’s private information will be, at least partially, revealed through market prices.
@@
With the above as motivation, we study how the insider’s awareness of price impact affects equilibrium price formulation and information transmission, as well as the indirect utility of both the insider and uninformed trader. We work in a single period normal-CARA model, seeking a linear price-impact equilibrium where a risk averse insider trades a bundle of risky assets with both a mass of uninformed risk-averse traders and liquidity providers, or noise traders. We differentiate the insider in two ways: first (clearly) in that she possesses private information; and second in that she accounts for the impact her trading activity has on prices. We then use this model to predict to what extent the insider’s signal is revealed to the market and how the uninformed traders correspondingly adjust their demands. We also compare equilibrium quantities of this model to those in the corresponding competitive economy model, to assess whether, and by how much, the informational content within equilibrium prices is reduced and what this implies in terms of traders’ demands.
We are particularly focused on the traders’ indirect utility, defined as the certainty equivalent from optimal trading. Our interest stems from a striking result in Grossman and Stiglitz (1980) (similar results are obtained in both Verrecchia (1982) and the recent Nezafat and Schroder (2023)) which shows that insider indirect utility need not be increasing in the quality of her signal. More precisely, if the signal is $G=X+Z_{I}$ where $X$ is the asset payoff and $Z_{I}$ is a noise term with precision $p_{I}$, then the map “$p_{I}\to\textrm{Insider Indirect Utility}$” takes one of the two forms shown in Figure 1.
Figure 1.
@@
This is problematic, because either it is never advantageous to obtain the signal (dot-dashed line) or it is only beneficial to obtain the signal for precisions below a certain threshold (solid line). As presumably there is a cost (in terms of money and/or effort) to produce the signal, the former case suggests price-taking equilibria with private signals are somewhat artificial, and the latter case requires the insider to estimate model parameters (in particular those related to other traders, which are difficult to estimate) to determine if it is worthwhile to refine her signal.
Motivated by Figure 1 we ask if our model produces a similar result, or if internalization of price impact ensures that the insider’s indirect utility is increasing in her signal precision. If this is indeed the case, it would suggest the improvement is driven by the insider’s differentiation in both access to better information and the internalization of price impact.
@@
We adjust the single period CARA-normal setting of Grossman and Stiglitz (1980) [^1] The competitive economy assumption in Grossman and Stiglitz (1980) is ubiquitous in the heterogeneous information asset pricing literature. It holds in the seminal papers of Grossman (1976), Grossman and Stiglitz (1980) and Hellwig (1980), and with the exception of the literature strand started by Kyle (1985); Back (1992), Rochet and Vila (1994) and Subrahmanyam (1991), price taking has remained the dominant assumption. by allowing the insider to internalize her price impact, while maintaining the presence of price-taking uninformed traders and liquidity providers [^2] Assuming the uninformed traders are price takers is realistic, as they represent a mass of small risk-averse traders who rationally optimize their positions, but do not have the power to move prices.. Following the related literature (e.g. Kyle (1985); Rochet and Vila (1994)), we study a linear impact equilibrium where the insider perceives the market price to be an affine function of the sum of her and the noise traders’ demand [^3] Linear price impact is common in the literature (see among others Kyle (1989), Vayanos (1999), Vives (2011)), and the affine structure of price impact is also seen (though not perceived a-priori by the insider) in the price-taking, or competitive, equilibrium (see (21) below)., with the affine coefficients endogenously determined through market clearing.
@@
In Theorem 1.2 we establish existence of a linear price impact (PI) equilibrium, where the coefficients are governed by the unique positive root to the cubic equation in (6) [^4] This stands in contrast to Subrahmanyam (1991) where equilibria are governed by solutions to a quintic equation. The difference arises because, consistent with Grossman and Stiglitz (1980) (see also Rochet and Vila (1994)) we assume the insider, by viewing both her demand and the resultant price, is able to deduce the noise trader’s demand.. To facilitate comparison, we summarize the price-taking (PT) equilibrium results of Grossman and Stiglitz (1980) in Proposition 1.3. Lastly, in both the PI and PT cases, we establish equilibria absent private information by showing it coincides with the zero precision limit (Proposition 1.4).
@@
Comparing the PI and PT equilibria, we first show, similarly to Kacperczyk et al. (2024), the public signal is fuzzier in the PI equilibrium (Proposition 2.1). This is reasonable, as the insider has a motive to hide her signal when submitting her demand, and hence her trading lowers the precision of the signal revealed by the prices. The uninformed traders duly recognize the insider reveals a muddied signal, and respond with a less elastic demand function. This in turn makes the PI equilibrium price less reactive to the public information than the PT price (consistent with the adverse-selection concerns of Kyle (1989) and Lou and Rahi (2023)). These results are robust across model parameter values, including agents’ risk aversion. We conclude that by assuming the insider is a price taker, one implicitly assumes the market receives a more precise signal, and prices are more reactive to public information, than when one assumes the insider internalizes price impact.
We next consider the traders’ indirect utility (henceforth referred to as “utility” for ease of exposition). Our main result shows that insider ex-ante utility is monotonically increasing in the private signal precision (Theorem 3.2). Therefore, the relationships displayed in Figure 1, which hold when either (i) all agents are price takers (e.g. Grossman and Stiglitz (1980)) or (ii) all agents internalize price impact (e.g. Nezafat and Schroder (2023)) do not occur in our model. This means there is always an incentive for the insider to refine her signal, provided she is differentiated from the other agents in terms of both (i) price impact internalization and (ii) asymmetric information.
It is of equal interest to understand the roles these two channels play in determining the insider utility in the first place. This question is motivated both by Figure 1 which indicates that information asymmetry is utility reducing, and by the naive belief that internalizing price impact should be utility improving. Thus, it is not immediately clear if on balance the PI equilibrium is utility improving over the PT equilibrium. Here, we first confirm that absent private information, internalizing price impact improves insider’s utility. Remarkably, this holds for any realization of the noise trader demand (Proposition 3.6). However, as surmised, the situation changes when one layers in the private signal. While typically (i.e. across the bulk of the parameter space) the utility benefit due to internalization of price impact outweighs the loss due to information asymmetry, and the PI model is utility improving over the PT model, this need not always be the case. Indeed, when the insider is sufficiently risk averse, the uninformed traders are sufficiently risk tolerant, and the variance of noise demand is sufficiently low [^5] The exact condition is given in equation (29) below. Also, the variance of noise traders’ demand is a proxy for the size of their trading, as discussed in Kovalenkov and Vives (2014); Nezafat and Schroder (2023)., the negative effects due to information asymmetry dominate and the PI model is utility reducing compared with the PT model. Therefore, there is no uniform utility ordering between the PI and PT equilibria, and in Section 4 we give further reasoning why this may occur, focusing on the insider’s demand function.
@@
We have shown that a better signal is utility improving for the insider in the PI model, but may not be utility improving for the insider in the PT model. This result underscores an important, and clarifying, fact. As long as the insider internalizes her price impact, the PI equilibrium price cannot be driven to the corresponding PT equilibrium price. This is due to the uninformed traders’ optimal demand. Despite being price takers themselves, when determining their optimal demand, they recognize that the insider internalizes her price impact. As price impact reduces the public signal’s precision, the uninformed traders demand is less elastic than in the PT equilibrium. The lower sensitivity to the public signal alters their demand function and hence, in the PI equilibrium the insider trades against a different residual demand than in the PT equilibrium. Thus, the PT equilibrium cannot be written as a special case of the PI one.
@@
Our paper contributes to the on-going literature on price-impact equilibria under asymmetric information; on market participants’ utility; and on the informativeness of equilibrium prices.
@@
In the PI equilibrium, the insider does not act as price-taker. Usually, the price-taking assumption is made for tractability, as in its absence one must specify a price impact model, and depending upon the specification, it may be very difficult to establish equilibria. In the aforementioned Kyle (1985); Back (1992); Rochet and Vila (1994), the insider’s demand is combined with exogenous noise traders’ demand before being sent to a risk neutral market maker who prices in a competitive environment. In Subrahmanyam (1991), market makers are risk averse and quote prices to remain at utility indifference (as opposed the uninformed agent of Grossman and Stiglitz (1980) who can be thought of as a market maker who quotes utility-optimal prices), but the insider does not know the noise trader demand before submitting her order (this is also the case in Kyle (1985)). In fact, our analysis updates Subrahmanyam (1991) in two directions. First, by assuming the uninformed trader is a utility optimizer, and second, by allowing the insider to the identify the noise trader demand through the public equilibrium price (as in Rochet and Vila (1994)).
@@
Following the seminal work of Kyle (1989), several models in the normal-CARA setting with price-impact and asymmetric information have been developed. For example, Vayanos (2001) and Rostek and Weretka (2015a) study dynamic thin markets with and without market makers respectively; Rostek and Weretka (2015b) focuses on the traders’ interdependent preferences and correlated private signals; and Malamud and Rostek (2017) and Anthropelos and Kardaras (2024) consider decentralized exchanges and restricted participation settings. Bergemann et al. (2021) studies a market of divisible goods where agents receive correlated signals, and their demand affects the revealed signal at equilibrium as in our model. In these works, as in Kyle (1989) and Vives (2011), strategic agents submit demand schedules which leads to a Nash equilibrium. In contrast to our model, therein all (non-noise) agents are strategic, and private information appears in the investors’ endowments. An overview of this literature is provided in Rostek and Yoon (2024).
@@
While large financial institutions invest to obtain private information, even when trading in markets that are not thin, theoretical studies such as ours that justify the positive relation between better information and higher gains from trading are scarce. In a competitive market, the fact that private information has positive value if it is internalized by an investor was pointed out in Hirshleifer (1971), with similar positive effects shown in a competitive model with a continuum of traders in Morris and Shin (2002). In non-competitive market settings, information acquisition has been studied in Vives (2011), Rostek and Weretka (2012) and Vives (2014). An extension of these papers (as well as Verrecchia (1982) which takes place in a competitive economy) to a two-stage model has been recently developed in Nezafat and Schroder (2023), where in the first stage one group of agents selects the precision of their private signal before trading. However, unlike our model, both the insider and uninformed traders are assumed strategic, which essentially makes the market thin. Strategic uninformed traders, together with specific conditions on noise traders’ demand, lead to an equilibrium at which private information is utility-deteriorating. By contrast, we show that the existence of zero-information equilibria are not possible if the uninformed traders are price takers, meaning that private information does have positive value for the insider.
@@
Kacperczyk et al. (2024) considers strategic informed and price-taking uninformed traders, as we do, where the role of initial endowments is highlighted for different types of informed traders. In Gong et al. (2022) the insider is assumed risk neutral and the role of uninformed traders is played by a market maker with a quadratic objective. A linear equilibrium, similar to Kyle (1985), is derived where the price under-reacts to public signal, as in our case. Strategic agents and asymmetric information have been included in Lou and Rahi (2023), where in a non-competitive market (in line with Kyle (1989) and Rostek and Weretka (2015a)) traders receive different ex-ante random values of a single asset (and potentially different private signals too). In contrast to our model (and to Kyle (1989)) there are no liquidity providers, which essentially implies that price informativeness does not change due to price impact. As in our model, there are conditions that lead to higher ex-ante expected utility for the uninformed traders than the insider. We reach a similar conclusion, but without assuming that uninformed traders act strategically.
@@
Lastly, another channel that reduces insider’s gains from trading due to private information is that of information sharing. For example, Goldstein et al. (2025) obtain this result in a model (based on Kyle (1989)) where informed traders share their private signals before trading (as in Indjejikian et al. (2014)).
@@
The rest of the paper is organized as follows. In Section 1, we provide the model and establish existence of the equilibria under consideration. Section 2 develops quantitative analysis and qualitative discussion on information transmission and signal and price sensitivities. Section 3 focuses on traders’ utility in different equilibria, and Section 4 examines equilibria structure regarding prices and risk allocation. Section 5 concludes with a discussion on the model’s predictions. An extension to a general multi-asset model is provided in Appendix A and all proofs are provided in Appendix B.